OpenStudy (anonymous):

On Tuesday, a local hamburger shop sold a combined total of 388 hamburgers and cheeseburgers. The number of cheeseburgers sold was three times the number of hamburgers sold. How many hamburgers were sold on Tuesday?

OpenStudy (anonymous):

h + c = 388 c = 3h now sub in 3h for c in the first equation

OpenStudy (anonymous):

4h=388 h=97

OpenStudy (anonymous):

yes..97 hamburgers.

OpenStudy (anonymous):

can you help me with another one?

OpenStudy (anonymous):

I will try

OpenStudy (anonymous):

A washer and a dryer cost $740 combined. The washer costs $90 more than the dryer. What is the cost of the dryer?

OpenStudy (anonymous):

w + d = 740 w = d + 90 sub d + 90 in for w in the 1st equation

OpenStudy (anonymous):

2d+90=740 2d=650 d=325

OpenStudy (anonymous):

you got it....dryer is 325

OpenStudy (anonymous):

okay I have one more question Alvin's age is three times Elga's age. The sum of their ages is 112 . What is Elga's age?

OpenStudy (anonymous):

a = 3e a + e = 112 sub in 3e for a in the 2nd equation

OpenStudy (anonymous):

4e=112 e=28

OpenStudy (anonymous):

correct :)

OpenStudy (anonymous):

thank you so much again!

OpenStudy (anonymous):

glad to help :)